# In the Queue: Understanding How Reddit Moderators Use the Modqueue

### v1

@@ 5.2. RQ1: How Does the Modqueue Fit into Moderators’ Moderation Workflows?
Overall, we found that respondents fit the modqueue into their workflows in pretty different ways. A small number avoided it almost entirely, relying instead on automation or lighter-weight in-thread actions (Section 5.2.1). Most used it as a kind of checklist, working through filtered and reported items as part of their everyday routine (Section 5.2.2. And a smaller set described using it more strategically, scanning for patterns or clusters of issues that signaled broader problems in the community (Section 5.2.3. In practice, this means the modqueue isn’t a single, uniform workspace, but one that takes on different roles depending on moderators’ philosophies, the kinds of communities they moderate, and the types and volumes of issues those communities face.
@@ 5.2.2. Everyday Uses: Reviewing Filtered Content, Responding to User Reports, and Clearing the Queue
Unsurprisingly, the most prominent use of the modqueue among the 106 moderators who reported using it was to review content automatically removed by platform-level filters or Automod. Several mods emphasized that this filtered content isn’t visible elsewhere, making the modqueue the only place where they can review and potentially release it to ensure that legitimate contributions aren’t mistakenly withheld from the community. Also expected, mods reported addressing user-submitted reports from the modqueue. Some moderators noted that this wasn’t their primary reason for using the interface, and described responding to user reports more opportunistically during regular subreddit perusal, or only after reviewing filtered content first.
@@ 5.2.3. Less Obvious Uses: The Modqueue as an Activity Radar or Moderation Hub
Beyond the more expected uses of the modqueue, a few mods described their approach in ways that suggest they use it as a kind of activity radar, scanning for patterns, emerging issues, or signs of disruption to get a sense of the community’s temperature or surface potential problems. After spotting notable trends in the modqueue (e.g. clusters of similar reports or contentious comment threads that seemed to need broader investigation or intervention), these mods reported leaving the modqueue interface to take more targeted or multi-step actions directly from the content on the subreddit page. As one respondent put it (when offering advice on how new mods should approach the modqueue):
@@ 5.3.1. Sequential or Priority-less Approaches
A little over half (56) of our respondents that reported using the modqueue described simple, sequential ways of moving through the modqueue as at least some part of the strategies they employ. Of these responses, many included mentions of traversing the modqueue in a particular direction, i.e. “top to bottom” or “bottom to top” (with no reference to how the modqueue may have been ordered. TThese responses often emphasized that what mattered most was addressing every item in the queue, regardless of order or prioritization. A handful of moderators echoed the view of P039, who explained that they had no specific strategy at all, seeing the goal simply as clearing the queue in its entirety: “All that is important to me is reviewing all reports. The order is irrelevant.”
Some moderators who described sequential traversal of the modqueue specifically mentioned working from either “newest first” or “oldest first.” Those who preferred newest-first approaches often framed this choice as a way to prioritize activity that was happening in real-time, suggesting that more recent reports warranted quicker attention. By contrast, some of the mods that processed oldest items first, like P038, indicated that their ordering preference was motivated by broader-values like fairness: “I deal with the oldest queries first, as that seems the fairer option.” Other reasons for preferring to go in order of oldest-first, specifically for filtered items, was to avoid leaving potentially valid posts in limbo.
@@ 5.3.3. Multi-Step or Adaptive Prioritization
I first search Posts in the modqueue. I scan through the list looking for obvious problems. I mentally prioritize posts that have gone live and have comments. Then I look at posts being held by Crowd Control [^2] Crowd Control is a moderation feature on Reddit that automatically filters or restricts comments from users based on their account status (e.g. if a user is very new or has very low karma) designed to automatically filter or restrict comments on posts based on the commenter’s account. or by the automod. (P081)
@@ 5.4.1. Reviewing the Broader Conversation Context
Mods also described using this opportunity to identify other posts or comments in the same thread that needed attention. In some cases, leaving the modqueue to investigate a single report led to broader moderation work, echoing earlier descriptions of the modqueue as an activity radar (LABEL:sec:findings:usage:activity):
@@ 6. Discussion
Our study surfaces the diverse, strategic, and sometimes improvised ways Reddit moderators use the modqueue to conduct mod review in practice, as well as the challenges they face while doing so. In this section, we discuss the main takeaways of our findings, outline implications for feature and platform design, and highlight limitations and directions for future work.
@@ 6.1.1. Human Oversight of Automated Decisions Still Relies on the Modqueue
6.1.2. Interfaces Shape Review Workflows, But Mods’ Values Shape Which Interfaces Are Used
@@ 6.1.3. Mods Leave the Modqueue to Gather Context Despite Built-In Features
Reviewing reports and making appropriate moderation decisions often requires information beyond what’s shown in the report itself. Reddit’s introduction of contextual panels in the New Reddit modqueue suggests an awareness of this, yet many mods still reported leaving the modqueue to gather necessary context. This includes checking surrounding conversations, viewing user histories, or looking up past mod actions—all of which are technically supported by contextual panels. That mods still leave the interface suggests either a lack of awareness, a lack of trust, or that New Reddit is not their default interface for deep review.
@@ 6.1.4. Fragmentation and Tool Incompatibility Can Disrupt Workflows
Most mods in our survey reported using more than one version of the modqueue, often supplemented by third-party tools. This mix of interfaces can cause problems. For instance, the activity indicator in New Reddit’s card view only works if everyone on the team is using New Reddit, which isn’t always the case. And while some mods prefer New Reddit, others stick with Old Reddit because it’s more compatible with long-used extensions. As Reddit updates its UI, compatibility with third-party tools can break, creating tension between interface updates and workflow reliability. While offering multiple options gives mods flexibility, the lack of coordination or integration across these options can cause friction.
@@ 6.2.2. Supporting Modular Tooling Across Interfaces and Subreddits
Subreddit-specific interface settings could also help mods adapt to these differences, and ensure that coordination tools (like activity indicators) function as intended. For instance, if one large subreddit wants all mods using the activity indicator, they should be able to standardize that experience, specifically for review of that subreddit’s content, to avoid can majorly inhibit mod review in those communities. This may push back on mods existing, personal workflows, but a balance must be struck between accommodating the needs of individual mods and the needs of mod teams and communities.
@@ 6.2.3. Providing Functionality for Pattern-Recognition
Some mods described using the modqueue not just as a place to process items, but as a way to scan for larger issues—like patterns in reports or spikes in problematic content. This repurposing reveals an underutilized strength of the modqueue: it centralizes signals from across the subreddit. When reports are timely and accurate, this can help mods spot disruptions that wouldn’t be obvious from isolated posts. Currently, this kind of pattern recognition isn’t directly supported. Mods rely on instinct and experience, rather than interface features, to detect trends. Some third-party tools support tagging or grouping, but, as previously mentioned, those tools can break when Reddit’s UI changes. This makes pattern-based use of the modqueue harder for less experienced mods to adopt.
Future design work could help by letting mods group, annotate, or track related reports more easily. More ambitious ideas include building interfaces or tabs that surface or visualize trends, or consolidate other moderation signals, such as post frequency, user behavior, or modmail alerts. These features could help mods respond proactively, rather than reactively. This kind of attention management may be even more impactful to broader moderation efficacy than just improving functionality for item-by-item review.
@@ 6.3.3. Lack of Interface-Specific Insight
Although we intentionally kept our survey questions high-level to accommodate multiple interface configurations, this choice limited our ability to analyze how specific modqueue versions (Old Reddit, New Reddit, mobile app) shape moderator behavior. Many respondents used multiple interfaces in parallel, and some may not have known which features were specific to which interface. This ambiguity makes it difficult to tie reported practices or frustrations to particular interface designs or features (unless features mentioned were exclusive to one interface over another). As previously mentioned, future work could address these limitations by complementing self-reported data with other data-driven methods or studies. For example, studies that isolate interface-specific effects or examine how mixed-interface teams coordinate could provide a more precise view of how design differences shape practice.
@@ 6.3.5. Opportunities to Optimize Report Review Workflows
Many moderators in our study described approaching report review in a systematic, checklist-like manner, working through items one at a time until the backlog was cleared. Future work could investigate how different approaches to item-by-item review influence both efficiency and quality of outcomes. This includes not only the strategies moderators already describe, such as prioritizing severe or time-sensitive reports, deferring ambiguous cases, or grouping similar items, but also the use of additional tools or interface features that support these strategies. Building a clearer understanding of how such approaches affect moderator performance could inform targeted improvements to existing workflows or motivate the development of new tools that better align with moderators’ needs. Such improvements would not replace the need for human judgment, but could reduce the routine effort involved in straightforward review and free up moderators’ attention for more complex or high-stakes cases.

### v2

# “Think about it like you’re a firefighter”: Understanding How Reddit Moderators Use the Modqueue

@@ 5.2. RQ1: How Does the Modqueue Fit into Moderators’ Moderation Workflows?
Overall, we found that respondents fit the modqueue into their workflows in pretty different ways. A small number avoided it almost entirely, relying instead on automation or lighter-weight in-thread actions (Section 5.2.1). Most used it as a kind of checklist, working through filtered and reported items as part of their everyday routine (Section 5.2.2). And a smaller set described using it more strategically, scanning for patterns or clusters of issues that signaled broader problems in the community (Section 5.2.3). In practice, this means the modqueue is not a single, uniform workspace, but one that takes on different roles depending on moderators’ philosophies, the kinds of communities they moderate, and the types and volumes of issues those communities face.
@@ 5.2.2. Everyday Uses: Reviewing Filtered Content, Responding to User Reports, and Clearing the Queue
Unsurprisingly, the most prominent use of the modqueue among the 106 moderators who reported using it was to review content automatically removed by platform-level filters or Automod. Several mods emphasized that this filtered content is not visible elsewhere, making the modqueue the only place where they can review and potentially release it to ensure that legitimate contributions are not mistakenly withheld from the community. Also expected, mods reported addressing user-submitted reports from the modqueue. Some moderators noted that this was not their primary reason for using the interface, and described responding to user reports more opportunistically during regular subreddit perusal, or only after reviewing filtered content first.
@@ 5.2.3. Less Obvious Uses: The Modqueue as an Activity Radar or Moderation Hub
Beyond the more expected uses of the modqueue, a few mods described their approach in ways that suggest they use it as a kind of activity radar, scanning for patterns, emerging issues, or signs of disruption to get a sense of the community’s temperature or surface potential problems. After spotting notable trends in the modqueue (e.g., clusters of similar reports or contentious comment threads that seemed to need broader investigation or intervention), these mods reported leaving the modqueue interface to take more targeted or multi-step actions directly from the content on the subreddit page. As one respondent put it (when offering advice on how new mods should approach the modqueue):
@@ 5.3.1. Sequential or Priority-less Approaches
A little over half (56) of our respondents that reported using the modqueue described simple, sequential ways of moving through the modqueue as at least some part of the strategies they employ. Of these responses, many included mentions of traversing the modqueue in a particular direction, i.e., “top to bottom” or “bottom to top” (with no reference to how the modqueue may have been ordered. These responses often emphasized that what mattered most was addressing every item in the queue, regardless of order or prioritization. A handful of moderators echoed the view of P039, who explained that they had no specific strategy at all, seeing the goal simply as clearing the queue in its entirety: “All that is important to me is reviewing all reports. The order is irrelevant.”
Some moderators who described sequential traversal of the modqueue specifically mentioned working from either “newest first” or “oldest first.” Those who preferred newest-first approaches often framed this choice as a way to prioritize activity that was happening in real-time, suggesting that more recent reports warranted quicker attention. By contrast, some of the mods that processed oldest items first, like P038, indicated that their ordering preference was motivated by broader-values like fairness: “I deal with the oldest queries first, as that seems the fairer option.” Other reasons for preferring an oldest-first order, specifically for filtered items, was to avoid leaving potentially valid posts “in limbo.”
@@ 5.3.3. Multi-Step or Adaptive Prioritization
I first search Posts in the modqueue. I scan through the list looking for obvious problems. I mentally prioritize posts that have gone live and have comments. Then I look at posts being held by Crowd Control [^2] Crowd Control is a moderation feature on Reddit that automatically filters or restricts comments from users based on their account status (e.g., if a user is very new or has very low karma) designed to automatically filter or restrict comments on posts based on the commenter’s account. or by the automod. (P081)
@@ 5.4.1. Reviewing the Broader Conversation Context
Mods also described using this opportunity to identify other posts or comments in the same thread that needed attention. In some cases, leaving the modqueue to investigate a single report led to broader moderation work, echoing earlier descriptions of the modqueue as an activity radar (Section 5.2.3):
@@ 6. Discussion
Our study surfaces the diverse, strategic, and sometimes improvised ways in which Reddit moderators use the modqueue to conduct mod review in practice, as well as the challenges they face while doing so. In this section, we discuss the main takeaways of our findings, discuss implications for feature and platform design, and highlight limitations and directions for future work.
@@ 6.1.1. Human Oversight of Automated Decisions Still Relies on the Modqueue
6.1.2. Interfaces Shape Review Workflows, But Mods’ Values Shape How Interfaces Are Used
@@ 6.1.3. Mods Leave the Modqueue to Gather Context Despite Built-In Features
Reviewing reports and making moderation decisions often requires information beyond what is shown in the report itself. Reddit’s introduction of contextual panels in the New Reddit modqueue suggests an awareness of this, yet many mods still reported leaving the modqueue to gather necessary context. This includes checking surrounding conversations, viewing user histories, or looking up past mod actions—all of which are technically supported by contextual panels. That mods still leave the interface suggests either a lack of awareness, a lack of trust, or that New Reddit is not their default interface for deep review.
@@ 6.1.4. Fragmentation and Tool Incompatibility Can Disrupt Workflows
Most mods in our survey reported using more than one version of the modqueue, often supplemented by third-party tools. This mix of interfaces can cause problems. For instance, the activity indicator in New Reddit’s card view only works if everyone on the team is using New Reddit, which is not always the case. And while some mods prefer New Reddit, others stick with Old Reddit because it is more compatible with long-used extensions. As Reddit updates its UI, compatibility with third-party tools can break, creating tension between interface updates and workflow reliability. While offering multiple options gives mods flexibility, the lack of coordination or integration across these options can cause friction.
@@ 6.2.2. Supporting Modular Tooling Across Interfaces and Subreddits
Subreddit-specific interface settings could also help mods adapt to these differences, and ensure that coordination tools (like activity indicators) function as intended. For instance, if one large subreddit wants all mods using the activity indicator, they should be able to standardize that experience, specifically for review of that subreddit’s content, to improve coordination among their mods. This may push back on mods existing, personal workflows, but a balance must be struck between accommodating the needs of individual mods and the needs of mod teams and communities.
@@ 6.2.3. Providing Functionality for Pattern-Recognition
Some mods described using the modqueue not just as a place to process items, but as a way to scan for larger issues—like patterns in reports or spikes in problematic content. This repurposing reveals an underutilized strength of the modqueue: it centralizes signals from across the subreddit. When reports are timely and accurate, this can help mods spot disruptions that wouldn’t be obvious from isolated posts. Currently, this kind of pattern recognition is not directly supported. Mods rely on instinct and experience, rather than interface features, to detect trends. Some third-party tools support tagging or grouping, but, as previously mentioned, those tools can break when Reddit’s UI changes. This makes pattern-based use of the modqueue harder for less experienced mods to adopt.
Future design work could help by letting mods group, annotate, or track related reports more easily. This could involve augmenting the modqueue with features tested in recent work on visualization tools to guide moderator attention and summarize trends/patterns in online activity (Choi et al., 2023; Liu et al., 2025). More ambitious ideas include building interfaces or tabs that surface or visualize trends, or consolidate other moderation signals, such as post frequency, user behavior, or modmail alerts. These features could help mods respond proactively, rather than reactively. This kind of attention management may be even more impactful to broader moderation efficacy than just improving functionality for item-by-item review.
@@ 6.3.3. Lack of Interface-Specific Insight
Although we intentionally kept our survey questions high-level to accommodate multiple interface configurations, this choice limited our ability to analyze how specific modqueue versions (e.g., Old Reddit, New Reddit, mobile app) shape moderator behavior. Many respondents used multiple interfaces in parallel, and some may not have known which features were specific to which interface. This ambiguity makes it difficult to tie reported practices or frustrations to particular interface designs or features (unless features mentioned were exclusive to one interface). As previously mentioned, future work could address these limitations by complementing self-reported data with other data-driven methods or studies. For example, studies that isolate interface-specific effects or examine how mixed-interface teams coordinate could provide a more precise view of how design differences shape practice.
@@ 6.3.5. Opportunities to Optimize Report Review Workflows
Many moderators in our study described approaching report review in a systematic, checklist-like manner, working through items one at a time until the backlog was cleared. Future work could investigate how different approaches to item-by-item review influence both efficiency and quality of outcomes. This includes not only the strategies moderators already describe—such as prioritizing severe or time-sensitive reports, deferring ambiguous cases, or grouping similar items—but also the use of additional tools or interface features that support these strategies. Building a clearer understanding of how such approaches affect moderator performance could inform targeted improvements to existing workflows or motivate the development of new tools that better align with moderators’ needs. Such improvements would not replace the need for human judgment, but could reduce the routine effort involved in straightforward review and free up moderators’ attention for more complex or high-stakes cases.